OpenStudy (anonymous):

Determine the midpoint given the following points A(4,2) B (-8, 18) 1) (2, -10) 2) (-2, 10) 3) (2, 10) 4) (-2, -10)

OpenStudy (campbell_st):

the midpoint is average of the x values and average of the y values \[x = \frac{4 - 8}{2}... y = \frac{2 + 18}{2}\] I'll let you don the calculation